A spring-mass system with m = 0.01 kg,
spring constant k = 36 Nm-1 and damping
constant y = 0.5 kg s-1 is subject to a
harmonic driving force. Calculate its
natural frequency and the resonance
frequency. How do they compare ? Discuss
the physical significance of your
result.
1
Expert's answer
2019-04-30T09:46:50-0400

2nd Newtons law:


"ma = -kx-yv \Rightarrow m\ddot{x}+y\dot{x}+kx=0"

Characteristic equation:


"m\lambda^2+y\lambda+k=0 \rightarrow \lambda = \frac{-y\pm i\sqrt{4 k m - y^2}}{2 m}"

Therefore, natural frequency


"\omega = \frac{\sqrt{4 k m - y^2}}{2 m}=\sqrt{ \frac{k}{m} - \frac{y^2}{4m^2}} = 54.5436\, \text{s}^{-1}"

Assume external force with amplitude A and frequency "\Omega". 2nd Newtons law takes the form


"m\ddot{x}+y\dot{x}+kx=A\sin{\Omega t}"

Solution:


"x(t)=A\frac{ \sin (t \Omega +\phi )}{\sqrt{(k-m\Omega^2)^2 + y^2\Omega^2}}+\text{homogeneous eq. solution}"

Resonance occurs at

"\Omega = \sqrt{k/m}=60\, \text{s}^{-1}"

Resonance frequency greater than natural freq. and doesn't depend on damping constant.